OpenStudy (anonymous):

+ 6 2 6 3 2 3 + square root of 2 plus square root of 6 divided by 3 answer please need help

OpenStudy (anonymous):

\[\sqrt{2}+\sqrt{6\div \sqrt{3}}\]